OpenStudy (anonymous):

Animals marking their territory with urine is an example of _______. a. competition b. predation c. commensalism d. mutualism

OpenStudy (anonymous):

Animals marking their territory with urine is an example of competition

OpenStudy (anonymous):

thanks @TwinYang

OpenStudy (abhisar):

By marking their territory they indicate or warn other animals to stay away from their territory. In this way they tackle competition.

OpenStudy (anonymous):

as always @Abhisar has a better response then me

OpenStudy (anonymous):

lol

OpenStudy (abhisar):

No it's ok @TwinYang ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (anonymous):

HaHa